OpenStudy (lilai3):

Please help! It will be greatly appreciated!!! Will give a medal and fan you! XD After a visit to the mall, Ray and Mary counted their money to see how much they had left. Ray said, "If I had $8 more, I would have as much as you." Mary replied, "If I had $8 more, I would have twice as much as you." Explain.

OpenStudy (anonymous):

R + 8 = M M + 8 = 2R

OpenStudy (anonymous):

system of equation now what is the question asking?

OpenStudy (anonymous):

\[R+8=2M\]\[M+8=2R\]So...\[R-2M=-8\]\[-2R+M=-8\]Then\[-2R+M+2(R-2M)=-8+2 \times -8\]\[-2R+M+2R-4M=-24\]\[-3M=-24\]\[M=8\] It's pretty clear that R=8 also.

OpenStudy (anonymous):

R = 16 M = 24

OpenStudy (anonymous):

Aylin got the first set wrong

OpenStudy (lilai3):

@Aylin i don't really get what you did, there....

OpenStudy (lilai3):

thank you, @Argos !!!

OpenStudy (anonymous):

Ray has 16 therefore if he had 8 more he would have the same as Mary 24 If Mary had 24 and got 8 more she would have 32 which is twice as much as 16Ray

OpenStudy (anonymous):

Ahh, I misread part of the question. Ignore my previous post please.